# Paper/main_online_appendix

### v3

@@ 2 Theory of Comparison Complexity
Let $X$ denote the set of options, and let $v_{x}$ denote the value of each $x\in X$. We consider a decision-maker (DM) who is uncertain about the value of each option, and when faced with a binary choice $\{x,y\}$, chooses based on a noisy signal on how $v_{x}$ and $v_{y}$ compare. In particular, the DM has continuous, i.i.d. priors over $v_{x}$ for all $x\in X$ distributed according to a symmetric distribution $Q$, and observes a signal $s_{xy}$ on the ordinal value comparison between $x$ and $y$, given by
@@
and chooses the option with the highest posterior expected value. Here, the precision $\tau_{xy}$ governs the ease of comparison between $x$ and $y$. Letting $\rho(x,y)$ denote the likelihood of choosing $x$ over $y$ conditional on the true values $v_{x}$ and $v_{y}$, we have $\rho(x,y)=\Phi(\text{sgn}(v_{x}-v_{y})\sqrt{\tau_{xy}})$, for $\Phi$ the standard normal CDF. [^1] In particular, $\rho(x,y)\equiv\mathbb{P}(\mathbb{E}[v_{x}|s_{xy}]>\mathbb{E}[v_{y}|s_{xy}]\,|\,v)$, where the DM randomizes if $\mathbb{E}[v_{x}|s_{xy}]=\mathbb{E}[v_{y}|s_{xy}]$. That is, the decision-maker’s likelihood of choosing the higher-valued option is increasing in the ease of comparison $\tau_{xy}$. Let $\tau:\{(x,y)\in X^{2}:x\neq y\}\to\mathbb{R}^{+}$ denote the associated mapping from each pair of options to their ease of comparison. In what follows, we propose a theory of how $\tau$ depends on the structure of choice options in the domains of multiattribute, lottery, and intertemporal choice.
@@ 2.1 Comparison Complexity: General Principles
First, we posit that holding fixed the value difference, options are easier to compare when they are more similar. Echoing work in psychology and economics (Tversky and Russo,, 1969; Rubinstein,, 1988; He and Natenzon, 2023a, ), this property reflects the intuition that if options are more similar, the DM can divert attention from the features that are similar across the options and so more easily assess the differences between them, thereby reducing the need to aggregate tradeoffs. Formally, we posit that the ease of comparison $\tau_{xy}$ is an increasing transformation $H$ of the value-dissimilarity ratio:
@@
Second, we posit that options are maximally easy to compare when they have a dominance relationship, i.e. when there are no tradeoffs. As formalized below, this principle gives rise to specific distance metrics given the domain-relevant dominance notion.
@@ 2.2 Multiattribute Choice
Consider the domain of multiattribute choice, where each option $x\in X\equiv X_{1}\times...\times X_{n}$ is defined on $n$ real-valued attributes, i.e. $X_{i}=\mathbb{R}$ [^2] In Appendix B.1, we provide guidance on how the analyst should specify the attributes in a given setting. Utility is linear in attributes, where the value of each option $x$ is given by $v_{x}=U(x)=\sum_{k}\beta_{k}x_{k}$ for attribute weights $\beta\in\mathbb{R}^{n}$. We propose that the ease of comparison is governed by the following representation:
@@
Note that this complexity representation satisfies the properties of similarity and dominance: holding fixed the value difference between two options, the ease of comparison is increasing in the similarity between $x$ and $y$ as measured by a distance metric, where this metric is chosen so that $\tau_{xy}$ achieves its maximal value in the presence of a dominance relationship. Specifically, if there is an attribute-wise dominance relationship between $x$ and $y$: i.e. $\beta_{k}x_{k}\geq\beta_{k}y_{k}$ for all $k$, the ease of comparison $\tau_{xy}$ takes on its maximal value of $H(1)$. [^3] To see this, note that by the triangle inequality we have $|\sum_{k}\beta_{k}(x_{k}-y_{k})|\leq\sum_{k}|\beta_{k}(x_{k}-y_{k})|$, and $|\sum_{k}\beta_{k}(x_{k}-y_{k})|=\sum_{k}|\beta_{k}(x_{k}-y_{k})|$ if and only if $x$ and $y$ have a dominance relationship.
@@ 2.2.1 Axiomatic Foundations
for some continuous, strictly increasing $G$ satisfying $G(r)=1-G(-r)$. [^5] There is a one-to-one correspondence between $G$ and $H$, which maps the value-dissimilarity ratio into signal precisions $\tau_{xy}$. In particular, for $r\in[0,1]$, $H(r)=(\Phi^{-1}(G(r))^{2}$, where $\Phi$ is the standard normal CDF.
Let $x_{\{k\}}y$ denote the option obtained by replacing the $k$th attribute of $y$ with $x_{k}$ [^6] That is $(x_{\{k\}}y)_{k}=x_{k}$ and $(x_{\{k\}}y)_{j}=y_{j}$ for all $j\neq k$.. Say that $x$ dominates $y$, written $x>_{D}y$, if $\rho(x_{\{k\}}y,y)\geq 1/2$ for all $k$ with a strict inequality for at least one $k$. Say that attribute $k$ is null if $\rho(x_{\{k\}}z,y_{\{k\}}z)=1/2$ for all $x,y,z\in X$. Consider the following conditions on $\rho$:
@@
Continuity and Linearity are standard axioms, and the latter reflects the fact that both utility and the $L_{1}$ distance are linear in attributes. Moderate Transitivity is a stochastic transitivity condition that allows choice probabilities to depend on a notion of the similarity between the options being compared, in addition to just their value difference. [^7] He and Natenzon (2023) show that in a finite domain, $\rho$ satisfies Moderate Transitivity if and only if $\rho(x,y)$ is increasing in the value difference between $x$ and $y$, normalized by some distance metric $d(x,y)$. While this result does not apply to our domain as it is not finite, our other axioms can be interpreted as adding structure to this distance metric.
Dominance and Simplification are the exact counterparts of the properties of $\tau^{L1}$ discussed earlier, stated in terms of choice probabilities. Dominance says that if $x$ is revealed better than $y$ on every attribute, then the likelihood of correctly choosing $x$ takes on its maximal value. Simplification says that eliminating the value difference between $x$ and $y$ along some attribute $i$ and redistributing that value to another attribute $j$ increases the DM’s likelihood of correctly choosing $x$ [^8] Hammond et al., (1998) argue on normative grounds that such an operation can be used to rationally simplify tradeoffs between multi-dimensional options.
@@
Suppose that all attributes are non-null and $n>2$. $\rho(x,y)$ has a $L_{1}$-complexity representation $(G,\beta)$ if and only if it satisfies M1–M5. Moreover, if $\rho(x,y)$ also has a $L_{1}$-complexity representation $(G^{\prime},\beta^{\prime})$, then $G^{\prime}=G$ and $\beta^{\prime}=C\beta$ for some $C>0$.
@@ 2.3.1 Lotteries
[…]
This measure provides a formal foundation for empirical work which has documented a tight connection between the CDF distance and choice rates. In particular, Enke and Shubatt, (2023) and Erev et al., (2008) show that the performance of lottery choice models markedly improves when choice noise is allowed to vary with a special case of $d_{CDF}$ in which $u$ is linear. Fishburn, (1978) proposes and axiomatizes a closely related model of binary choice over lotteries in which choice probabilities are increasing in the CDF ratio. [^10] We thank Paulo Natenzon for bringing our attention to this paper. As discussed in Section 2.3.4 and Appendix B.1, we provide an alternative axiomatic foundation for this choice model using properties analogous to our multiattribute axioms. We also extend each of our models to multinomial choice.
@@ 2.4 Parameterizing the Model
In each domain, choice probabilities under our model take the form
$\rho(x,y)$ $=G\left(\frac{U(x)-U(y)}{d(x,y)}\right),$
where the signed value-dissimilarity ratio $\left(\frac{U(x)-U(y)}{d(x,y)}\right)$ is specified according to Definitions 1, 3, and 4, and $G$ is an increasing transformation satisfying $G(r)=1-G(-r)$. To obtain quantitative predictions, the analyst needs to specify the preference parameters that enter the value-dissimilarity ratio — the attribute weights $\beta$ in multiattribute choice, the Bernoulli utility $u$ in lottery choice, and the discount factor $\delta$ in intertemporal choice — as well as the transformation $G$. Each of these objects can be identified from binary choice data, as stated in Theorem 1 for multiattribute choice, and in Appendix B.1 for our other two domains.
@@ 3.1 Multinomial Choice Extension
Consider the same setting as in our binary choice framework. There is a set of options $X$, and the DM has continuous, i.i.d. priors over $v_{z}$ for all $z\in X$, distributed according to a symmetric distribution $Q$. Let $\mathcal{M}$ denote the collection of finite subsets of $X$, and let $\mathcal{A}=\{A\in\mathcal{M}:|A|\geq 2\}$ denote the set of finite menus. The DM faces a choice problem $(A,C)\in\mathcal{A}\times\mathcal{M}$, comprised of a menu of options $A$ and a choice context $C$ – a set of options the DM observes but cannot choose, i.e. phantom options. The DM chooses from $A$ based on signals on how each pair of options in $A\cup C$ compares.
@@ 3.2 Model Properties
This model has two key implications. First, it predicts that comparison complexity leads to noisy but unbiased choice in binary menus — the DM may err, but is more likely to pick the higher-value option. Second, it predicts systematic biases in choice from larger menus: when facing hard-to-compare alternatives, the DM relies on information from comparisons to other options, which can distort choice. This generates a) context effects when options are hard to compare to each other but differ in their ease of comparison to other options in the choice context, and b) systematic pull-to-center distortions in valuations as a function of how hard the option being valued is to compare to the numeraire.
@@ 3.2.1 Context Effects
When combined with our theory of comparison complexity in multiattribute choice, this result rationalizes familiar decoy and asymmetric dominance effects.
@@ 3.2.2 Compression Effects in Valuation
Here, $\rho((x^{1},...x^{n}),(A^{1},...,A^{n})|C)$ records the frequency of choosing $x^{i}\in A^{i}$ for $i=1,...,n$. Given an option $x$ and a price list $Z$, a valuation task $(x,Z)$ is simply the binary menu sequence $A^{1},...,A^{n}=\{x,z^{1}\},...,\{x,z^{n}\}$.
Since the DM learns the ranking of prices in $Z$, this procedure yields a single switching point: for any signal realization, there is an index $R\in\{1,...,n,n+1\}$ for which the DM chooses the option $x\in A^{k}$ for all $k\geq R$, and the price $z^{k}\in A^{k}$ for all $k<R$. $R$ reveals where the DM believes the object $x$ falls within the ranking of prices, i.e. the subject’s valuation in terms of $Z$. We will be interested in the distribution over $R$ induced by $\rho((x^{1},...x^{n}),(A^{1},...,A^{n})|Z)$, which we denote by $R(x,Z)$. [^15] Given a signal $s$, the DM’s posterior switching point $R$ is computed by calculating $\operatorname{\mathbb{E}}[v_{x}|s]$ and $\operatorname{\mathbb{E}}[v_{j}|s]$ for all $j\in\{1,2,...,n\}$, and finding the unique index $R$ such that $\operatorname{\mathbb{E}}[v_{x}|s]<\operatorname{\mathbb{E}}[v_{R-1}|s]$ and $\operatorname{\mathbb{E}}[v_{x}|s]>\operatorname{\mathbb{E}}[v_{R}|s]$ (in the case of ties, we assume the DM randomizes as described in Appendix B.3). With a slight abuse of notation, we denote $v_{k}\equiv v_{z^{k}}$ and $\tau_{xk}\equiv\tau_{xz^{k}}$.
Our model predicts that when $x$ is hard to compare to prices, valuations will be systematically compressed towards the center of the price list. Let $R^{*}(x,z)\in\{1,...,n,n+1\}$ denote the true ranking of $x$ relative to $Z$ [^16] That is, $v_{x}>v_{k}$ if $k\geq R^{*}(x,Z)$ and $v_{x}<v_{k}$ otherwise. For ease of exposition, we assume $x$ is not indifferent to any price in $Z$.
@@
Given a valuation task $(x,Z)$, where $\tau_{xk}=\tau$ and $v_{x}\neq v_{k}$ for all $k=1,...,n$, we have the following:
@@ 3.3 Preference Reversals
$x:\quad\$23.50\text{ with 19\%}$%
$y:\quad\$4.75\text{ with 94\%}$%
Most subjects choose $y$ over $x$ in direct choice, yet state a higher certainty equivalent for $x$. A number of explanations for these reversals have been put forth, such as intransitive preferences or independence violations (see Seidl, (2002) for a review). Our model rationalizes these reversals as the consequence of the differential ease of comparing options to money. Under our complexity notion, $y$ is easier to compare to money than $x$, which results in differential pull-to-center distortions. As this force distorts valuations but not direct choice, reversals can occur. As we show below, this logic not only generates reversals in lottery choice: it predicts that similar reversals occur in intertemporal choice, and that these reversals can be eliminated by manipulating the ease of comparing each option to prices.

### v4

@@ 2 Theory of Comparison Complexity
Let $X$ denote the set of options, and let $v_{x}$ denote the value of each $x\in X$. We consider a decision-maker (DM) who is uncertain about the value of each option, and when faced with a binary choice $\{x,y\}$, chooses based on a noisy signal on how $v_{x}$ and $v_{y}$ compare. In particular, the DM has continuous, i.i.d. priors over each $(v_{x})_{x\in X}$ given by the symmetric distribution $Q$, and receives a signal $s_{xy}$ on the ordinal value comparison between $x$ and $y$:
@@
and chooses the option with the highest posterior expected value. [^1] Normality of $\epsilon_{xy}$ can be relaxed. Appendix C.2 shows that our results hold for any $\epsilon_{xy}$ with a continuous density that is symmetric around 0 and satisfies a monotone likelihood ratio property. Here, the precision $\tau_{xy}$ governs the ease of comparison between $x$ and $y$. Letting $\rho(x,y)$ denote the likelihood of choosing $x$ over $y$ conditional on the true values $v_{x}$ and $v_{y}$, we have $\rho(x,y)=\Phi(\text{sgn}(v_{x}-v_{y})\sqrt{\tau_{xy}})$, for $\Phi$ the standard normal CDF. [^2] In particular, $\rho(x,y)\equiv\mathbb{P}(\mathbb{E}[v_{x}|s_{xy}]>\mathbb{E}[v_{y}|s_{xy}]\,|\,v)$, where the DM randomizes if $\mathbb{E}[v_{x}|s_{xy}]=\mathbb{E}[v_{y}|s_{xy}]$. That is, the DM’s likelihood of choosing the higher-valued option is increasing in the ease of comparison $\tau_{xy}$. Let $\tau:\{(x,y)\in X^{2}:x\neq y\}\to\mathbb{R}^{+}$ denote the associated mapping from each pair of options to their ease of comparison. In what follows, we propose a theory of how $\tau$ depends on the structure of choice options in the domains of multiattribute, lottery, and intertemporal choice.
@@ 2.1 Comparison Complexity: General Principles
First, we posit that holding fixed the value difference, options are easier to compare when they are more similar. Echoing work in psychology and economics (Tversky and Russo,, 1969; Rubinstein,, 1988; He and Natenzon,, 2024), this property reflects the intuition that if options are more similar, the DM can divert attention from the features that are similar across the options and so more easily assess the differences between them, thereby reducing the need to aggregate tradeoffs. Formally, we posit that the ease of comparison $\tau_{xy}$ is an increasing transformation $H$ of the value-dissimilarity ratio:
@@
Second, we posit that options are maximally easy to compare when there are no tradeoffs – that is, when they have a dominance relationship. As formalized below, this principle gives rise to specific distance metrics given the domain-relevant dominance notion.
@@ 2.2 Multiattribute Choice
Consider the domain of multiattribute choice, where each option $x\in X\equiv X_{1}\times...\times X_{n}$ is defined on $n$ real-valued attributes, i.e. $X_{i}=\mathbb{R}$. [^3] In Appendix B.1, we provide guidance on how the analyst should specify the attributes in a given setting: specifically, how attributes can be identified from choices. Utility is linear in attributes, where the value of each option $x$ is given by $v_{x}=U(x)=\sum_{k}\beta_{k}x_{k}$ for attribute weights $\beta\in\mathbb{R}^{n}$. We propose that the ease of comparison is governed by the following representation:
@@
Note that this complexity representation satisfies the properties of similarity and dominance: holding fixed the value difference between two options, the ease of comparison is increasing in the similarity between $x$ and $y$ as measured by a distance metric, where this metric is chosen so that $\tau_{xy}$ achieves its maximal value in the presence of a dominance relationship. Specifically, if there is an attribute-wise dominance relationship between $x$ and $y$: i.e. $\beta_{k}x_{k}\geq\beta_{k}y_{k}$ for all $k$, the ease of comparison $\tau_{xy}$ takes on its maximal value of $H(1)$.
@@ 2.2.1 Axiomatic Foundations
for some continuous, strictly increasing $G$ satisfying G(r)=1G(r)G(r)=1-G(-r). [^5] There is a one-to-one correspondence between $G$ and $H$, which maps the value-dissimilarity ratio into signal precisions $\tau_{xy}$. In particular, for $r\in[0,1]$, $H(r)=(\Phi^{-1}(G(r))^{2}$, where $\Phi$ is the standard normal CDF.
Let $x_{\{k\}}y$ denote the option obtained by replacing the $k$th attribute of $y$ with $x_{k}$ [^6] That is $(x_{\{k\}}y)_{k}=x_{k}$ and $(x_{\{k\}}y)_{j}=y_{j}$ for all $j\neq k$.. Say that $x$ dominates $y$, written $x>_{D}y$, if $\rho(x_{\{k\}}y,y)\geq 1/2$ for all $k$ with a strict inequality for at least one $k$ – that is, if $x$ is revealed better than $y$ along each attribute when considering attributes in isolation. Say that attribute $k$ is null if $\rho(x_{\{k\}}z,y_{\{k\}}z)=1/2$ for all $x,y,z\in X$. Consider the following conditions on $\rho$:
@@
Continuity and Linearity are standard axioms, and the latter reflects the fact that both utility and the $L_{1}$ distance are linear in attributes. Moderate Transitivity is a stochastic transitivity condition that allows choice probabilities to depend on a notion of the similarity between the options being compared, in addition to just their value difference. [^7] He and Natenzon, (2024) show that in a finite domain, $\rho$ satisfies Moderate Transitivity if and only if $\rho(x,y)$ is increasing in the value difference between $x$ and $y$, normalized by some distance metric $d(x,y)$. While this result does not apply to our domain as it is not finite, our other axioms can be interpreted as adding structure to this distance metric.
Dominance and Simplification are the exact counterparts of the properties of $\tau^{L1}$ discussed earlier, stated in terms of choice probabilities. Dominance says that if $x$ is revealed better than $y$ on every attribute, then the likelihood of correctly choosing $x$ takes on its maximal value. Simplification says that eliminating the value difference between $x$ and $y$ along some attribute $i$ and redistributing that value to another attribute $j$ increases the DM’s likelihood of correctly choosing xx. [^8] Hammond et al., (1998) argue on normative grounds that such an operation can be used to rationally simplify tradeoffs between multi-dimensional options.
@@
Suppose that all attributes are non-null and $n>2$. $\rho(x,y)$ has an $L_{1}$-complexity representation $(G,\beta)$ if and only if it satisfies M1–M5. Moreover, if $\rho(x,y)$ also has an $L_{1}$-complexity representation $(G^{\prime},\beta^{\prime})$, then $G^{\prime}=G$ and $\beta^{\prime}=C\beta$ for some $C>0$.
@@ 2.3.1 Lotteries
This measure provides a formal foundation for empirical work documenting a tight connection between the CDF distance and choice rates (Enke and Shubatt,, 2023; Erev et al.,, 2008; Buschena and Zilberman,, 2008). In particular, this work shows that the performance of lottery choice models markedly improves when choice noise is allowed to vary with a special case of $d_{CDF}$ in which $u$ is linear. Fishburn, (1978) proposes and axiomatizes a closely related model of binary choice over lotteries in which choice probabilities are increasing in the CDF ratio. [^10] We thank Paulo Natenzon for bringing our attention to this paper. As discussed below and in Appendix B.1, we provide an alternative axiomatic foundation for this choice model using easily understood conditions analogous to our multiattribute axioms.
@@ 2.4 Parameterizing the Model
In each domain, choice probabilities under our model take the form $\rho(x,y)=G\left(\frac{U(x)-U(y)}{d(x,y)}\right)$, where the signed value-dissimilarity ratio $\left(\frac{U(x)-U(y)}{d(x,y)}\right)$ is specified according to Definitions 1, 3, and 4, and $G$ is an increasing transformation satisfying G(r)=1G(r)G(r)=1-G(-r). To obtain quantitative predictions, the analyst needs to specify the preference parameters that enter the value-dissimilarity ratio — the attribute weights $\beta$ in multiattribute choice, the Bernoulli utility $u$ in lottery choice, and the discount factor $\delta$ in intertemporal choice — as well as the transformation $G$. Each of these objects can be identified from binary choice data, as stated in Theorem 1 for multiattribute choice, and in Appendix B.1 for our other two domains.
@@ 3.1 Multinomial Choice Extension
Consider the same setting as in our binary choice framework. There is a set of options $X$, and the DM has continuous, i.i.d. priors over $v_{z}$ for all $z\in X$, distributed according to a symmetric distribution $Q$. Let $\mathcal{A}$ denote the collection of non-empty finite subsets of $X$, and let $\mathcal{C}=\mathcal{A}\cup\{\varnothing\}$. The DM faces a choice problem $(A,C)\in\mathcal{A}\times\mathcal{C}$, comprised of a menu of options $A$ and a choice context $C$ – a set of options the DM observes but cannot choose, i.e. phantom options. The DM chooses from $A$ based on signals on how each pair of options in $A\cup C$ compares.
@@ 3.2 Model Properties
This model has two key implications. First, it predicts that comparison complexity leads to noisy but unbiased choice in binary menus — the DM may err, but is more likely to pick the higher-value option. Second, it predicts systematic biases in choice from larger menus: when facing hard-to-compare alternatives, the DM relies on information from comparisons to other options, which can distort choice. This generates a) context effects when options are hard to compare to each other but differ in their ease of comparison to other options in the choice context, and b) systematic pull-to-center distortions in valuations as a function of the difficulty of comparing the option being valued to the numeraire.
@@ 3.2.1 Context Effects
When combined with our theory of complexity in multiattribute choice, this result rationalizes documented decoy and asymmetric dominance effects (e.g. Huber et al.,, 1982).
@@ 3.2.2 Compression Effects in Valuation
Here, $\rho((x^{1},...x^{n}),(A^{1},...,A^{n})|C)$ records the frequency of choosing $x^{i}\in A^{i}$ for i=1,,ni=1,...,n. Given an option $x$ and a price list $Z$, a valuation task $(x,Z)$ is simply the binary menu sequence $A^{1},...,A^{n}=\{x,z^{1}\},...,\{x,z^{n}\}$.
Since the DM learns the ranking of prices in $Z$, this procedure yields a single switching point: for any signal realization, there is an index $R\in\{1,...,n,n+1\}$ for which the DM chooses the option $x\in A^{k}$ for all $k\geq R$, and the price $z^{k}\in A^{k}$ for all $k<R$. $R$ reveals where the DM believes the object $x$ falls within the ranking of prices, i.e. the subject’s valuation in terms of $Z$. We will be interested in the distribution over $R$ induced by $\rho((x^{1},...x^{n}),(A^{1},...,A^{n})|Z)$, which we denote by R(x,Z)R(x,Z). [^15] Given a signal $s$, the DM’s posterior switching point $R$ is computed by calculating $\operatorname{\mathbb{E}}[v_{x}|s]$ and $\operatorname{\mathbb{E}}[v_{j}|s]$ for all $j\in\{1,2,...,n\}$, and finding the unique index $R$ such that $\operatorname{\mathbb{E}}[v_{x}|s]<\operatorname{\mathbb{E}}[v_{R-1}|s]$ and $\operatorname{\mathbb{E}}[v_{x}|s]>\operatorname{\mathbb{E}}[v_{R}|s]$ (in the case of ties, we assume the DM randomizes as described in Appendix B.3). With a slight abuse of notation, we denote $v_{k}\equiv v_{z^{k}}$ and $\tau_{xk}\equiv\tau_{xz^{k}}$.
Our model predicts that when $x$ is hard to compare to prices, valuations will be systematically compressed towards the center of the price list. Let $R^{*}(x,Z)\in\{1,...,n,n+1\}$ denote the true ranking of $x$ relative to ZZ. [^16] That is, $v_{x}>v_{k}$ if $k\geq R^{*}(x,Z)$ and $v_{x}<v_{k}$ otherwise. For ease of exposition, we assume $x$ is not indifferent to any price in $Z$.
@@
Given a valuation task $(x,Z)$, where $\tau_{xk}=\tau$ and $v_{x}\neq v_{k}$ for all k=1,,nk=1,...,n, we have the following:
@@ 3.3 Preference Reversals
$x:\quad\$23.50\text{ with 19\%}$
$y:\quad\$4.75\text{ with 94\%}$
Most subjects choose $y$ over $x$ in direct choice, yet state a higher certainty equivalent for $x$. A number of explanations for these reversals have been proposed, such as intransitive preferences or independence violations (see Seidl, (2002) for a review). Our model can rationalize these reversals between choice and valuation as expressions of the same underlying preferences, taking into account errors produced by the differential ease of comparing options to money. Under our complexity notion, $y$ is easier to compare to money than $x$, which results in differential pull-to-center distortions. As this force distorts valuations but not direct choice, reversals can occur. As we show below, this logic not only generates reversals in lottery choice: it predicts similar reversals in intertemporal choice, and that these reversals can be eliminated by manipulating the ease of comparing options to prices.